### v3

@@ 5.3 Influence of curved microchannel on the dynamic contact angle
Plotting the dynamic CA with calculated using local CL velocities, the comparison between experimental results and fit to the MKT model can be seen in fig. 13. The free parameters and estimated from curve-fitting are listed in table 5. It should be noted that the identical static CA ° is employed for all parts of variation 3, except for the inside of position 1, where ° is applied. That is attributed to the observed significantly small dynamic CAs, resulting from the formation of a distinct tip between the interface and the wall as presented in fig. 12(b). This strong bending of interface along the inside CL is mostly likely due to the considerable difference of the local CL velocities between the inner and the outer radius of the microchannel, causing the pinning of the inner radius CL. The different shear rates on curved region might also contribute to this phenomenon. In addition, it is worth stressing that the rectangular geometry of a channel is known to cause the “finger” effect of the CL along the sharp edges of channel ([10]; [34]), which might be enhanced by the pronounced curvature of variation 3.
@@
| Channel part | (deg) | () | () |
| Straight | 86 | 1.07 | 284.86 |
| Inside position 1 | 70 | 1.03 | 424.12 |
| Outside position 1 | 86 | 1.03 | 400.75 |
| Inside position 2 | 86 | 1.20 | 195.86 |
| Outside position 2 | 86 | 1.02 | 410.16 |

### v4

@@ 5.3 Influence of curved microchannel on the dynamic contact angle
Plotting the dynamic CA with calculated using local CL velocities, the comparison between experimental results and fit to the MKT model can be seen in fig. 14. The free parameters and estimated from curve-fitting with standard deviations are listed in table 7. It should be noted that the identical static CA ° is employed for all parts of variation 3, except for the inside of position 1, where ° is applied. That is attributed to the observed significantly small dynamic CAs, resulting from the formation of a distinct tip between the interface and the wall as presented in fig. 13(b). This strong bending of interface along the inside CL is mostly likely due to the considerable difference of the local CL velocities between the inner and the outer radius of the microchannel, causing the pinning of the inner radius CL. The different shear rates on curved region might also contribute to this phenomenon. In addition, it is worth stressing that the rectangular geometry of a channel is known to cause the “finger” effect of the CL along the sharp edges of channel ([16]; [49]), which might be enhanced by the pronounced curvature of variation 3.
@@
| Channel part | |||
| (deg) | () | () | |
| Straight | 86 | 1.03 0.05 | 385.61 124.52 |
| Inside1 | 70 | 1.03 0.11 | 424.12 153.90 |
| Outside1 | 86 | 1.02 0.03 | 457.26 90.52 |
| Inside2 | 86 | 1.34 0.09 | 107.37 43.47 |
| Outside2 | 86 | 1.03 0.04 | 406.92 111.90 |